CERTAIN APPLICATIONS OF THE BURNSIDE RINGS AND GHOST RINGS IN THE REPRESENTATION THEORY OF FINITE

GROUPS (II)

K. K. NWABUEZE F. VAN OYSTAEYEN

Abstract. Using the Burnside ring theoretic methods a new setting and a complete description of the Artin exponent A(G) of finite p-groups was obtained in a previous article of the first-named author. In this paper, we compute A(G) for any finite group G – hence providing the global version of that result.